OpenStudy (anonymous):

SOLVE: write an equation of the line ha passess through the given piont and is parallel to the graph of the given equation 1. (3,2);y = 3x - 2 2. (-8,6);y = -1/4x (fraction)+5

OpenStudy (slaaibak):

You already posted this and we did give you steps on how do do this

OpenStudy (anonymous):

parallel is the same slope

OpenStudy (anonymous):

do you know point slope formula?

OpenStudy (anonymous):

yes but i cant do it i tryed

OpenStudy (karatechopper):

show me what u tried

OpenStudy (anonymous):

i tryed doing it like he person said but i dino come out

OpenStudy (karatechopper):

u said u tried, show me the work u did

OpenStudy (anonymous):

i put y = 3x - 2 then i passed the y to the (3,2) then i got y = 3x +3y x 2y

OpenStudy (karatechopper):

o boy

OpenStudy (karatechopper):

do u know what point slope form is?

OpenStudy (anonymous):

what see im confuzzed!

OpenStudy (karatechopper):

lol u said u knew what it was!!

OpenStudy (anonymous):

piont slope form is y=mx+b rite

OpenStudy (karatechopper):

no tht is slope intercept form

OpenStudy (anonymous):

ok then its Ax+By=C

OpenStudy (anonymous):

rite?

OpenStudy (karatechopper):

no thts standard form

OpenStudy (anonymous):

omg i so confuzzed! can u help me

OpenStudy (karatechopper):

im here

OpenStudy (karatechopper):

y-y1=m(x-x1)

OpenStudy (karatechopper):

tht is point slope form

OpenStudy (anonymous):

ok so its y-2=3x(x-3)

OpenStudy (karatechopper):

well 3x is not your slope. 3 is your slope

OpenStudy (anonymous):

so its y-2=3(x-3)

OpenStudy (karatechopper):

correct to

OpenStudy (karatechopper):

now simplify tht

OpenStudy (anonymous):

simplfy?

OpenStudy (karatechopper):

maybe i should say solve for y now

OpenStudy (anonymous):

but i dnt have to simplfiy rite and the other one is .

OpenStudy (karatechopper):

yes u do have to simplify

OpenStudy (anonymous):

ugh ok so you _m each side byx

OpenStudy (karatechopper):

nooo

OpenStudy (anonymous):

then waht do i fdo because i have to geo off n a minute to go to bed im oly 12 and my dad get mad if i dn go to bed when he wants me to

OpenStudy (anonymous):

omg for real lol

OpenStudy (karatechopper):

yes. y=3x-7 <- theres ur answer now go to bed kiddo